50.0 ml of 2.50M Sulfuric acid reacted with 43.5ml 2.75 Aluminum hydroxide. The reaction yielded Aluminum sulfate and water. (Al=27g/mol, S=32g/mol, H=1g/mol, O=16g/mol)

1. Write the balance chemical equation.

2. What is the limiting reactant?

3. How much reactant is in excess?

4. How many grams of Aluminum sulfate will be produced?


1
Expert's answer
2022-04-19T03:11:54-0400

Solution:

Molar mass of sulfuric acid, H2SO4 = 2×Ar(H) + Ar(S) + 4×Ar(O) = 2×1 + 32 + 4×16 = 98 (g/mol)

The molar mass of H2SO4 is 98 g/mol

Molar mass of aluminum hydroxide, Al(OH)3 = Ar(Al) + 3×Ar(O) + 3×Ar(H) = 27 + 3×16 + 3×1 = 78 (g/mol)

The molar mass of Al(OH)3 is 78 g/mol

Molar mass of aluminum sulfate, Al2(SO4)3 = 2×Ar(Al) + 3×Ar(S) + 12×Ar(O) = 2×27 + 3×32 + 12×16 = 342 (g/mol)

The molar mass of Al2(SO4)3 is 342 g/mol


Calculate moles of each reactant:

Moles of H2SO4 = (2.50 M) × (50.0 mL) × (1 L / 1000 mL) = 0.125 mol H2SO4

Moles of Al(OH)3 = (2.75 M) × (43.5 mL) × (1 L / 1000 mL) = 0.1196 mol Al(OH)3


Balanced chemical equation:

3H2SO4(aq) + 2Al(OH)3(s) → Al2(SO4)3(aq) + 6H2O(l)


According to stoichiometry:

3 mol of H2SO4 react with 2 mol of Al(OH)3

Thus, 0.125 mol of H2SO4 react with:

(0.125 mol H2SO4) × (2 mol Al(OH)3 / 3 mol H2SO4) = 0.0833 mol Al(OH)3

However, initially there is 0.1196 mol of Al(OH)3 (according to the task)

Therefore, H2SO4 acts as limiting reactant and Al(OH)3 is excess reactant


Moles of Al(OH)3 in excess = 0.1196 mol − 0.0833 mol = 0.0363 mol

Moles of Al(OH)3 in excess = 0.0363 mol

Mass of Al(OH)3 in excess = (0.0363 mol) × (78 g/mol) = 2.83 g

Mass of Al(OH)3 in excess = 2.83 g


According to stoichiometry:

3 mol of H2SO4 produce 1 mol of Al2(SO4)3

Thus, 0.125 mol of H2SO4 produce:

(0.125 mol H2SO4) × (1 mol Al2(SO4)3 / 3 mol H2SO4) = 0.0417 mol Al2(SO4)3

Moles of Al2(SO4)3 = 0.0417 mol


The molar mass of Al2(SO4)3 is 342 g/mol

Therefore,

Mass of Al2(SO4)3 = (0.0417 mol) × (342 g/mol) = 14.26 g

Mass of Al2(SO4)3 = 14.26 g


Answers:

1) Balanced chemical equation: 3H2SO4(aq) + 2Al(OH)3(s) → Al2(SO4)3(aq) + 6H2O(l)

2) The limiting reactant is H2SO4

3) Moles of Al(OH)3 in excess = 0.0363 mol; Mass of Al(OH)3 in excess = 2.83 g

4) 14.26 grams of aluminum sulfate will be produced